OpenStudy (loser66):

Give an example of a non-zero alternating 2-linear form w on 3-dimensional space such that w(x1, x2) =0 for some set of linearly independent vector x1,x2 Please, help

OpenStudy (loser66):

@hartnn @Kainui @kirbykirby

OpenStudy (kirbykirby):

Hm I am not sure what an alternating 2-linear form is :S

OpenStudy (loser66):

bilinear.

OpenStudy (kirbykirby):

Hm I'm sorry my brain is too tired right now for linear algebra :( I can't think clearly .

OpenStudy (anonymous):

a better question is why are you doing homework on a friday night

OpenStudy (loser66):

@bh6180 I have a bunch of homework due on Tuesday. Although I started my homework right after the lecture and in weekend, I rarely have it done before due date. Friday?? what is special?

OpenStudy (loser66):

@amistre64

OpenStudy (amistre64):

does w: (x1,x2) to 0, for all x1,x2 in the set if linearlly independant vectors? if so, isnt that a kernel maybe? these are just some guesses that i have no clue about

OpenStudy (loser66):

@ikram002p I need understand the stuff . I am ok if the stuff is all letters, like w (x1,x2) = u (x2,x3) v (x1) +..... but really lost when apply to particular case like this

OpenStudy (loser66):

i have the solution here, but don't understand the logic on it.

OpenStudy (loser66):

My particular problem is n = 3 and k =2.

OpenStudy (ikram002p):

ohk , ill review it :O

OpenStudy (ikram002p):

humm i checked the txt book :O but sry have no clue abt it , seems i forget it :'(

OpenStudy (loser66):

It's ok, friend. I appreciate any trying. :)

OpenStudy (experimentx):

it seems the example is given in the solution ... what do you want more?

OpenStudy (loser66):

I want to know why the last column of det is (1,1,0)

OpenStudy (loser66):

and how to defined \(Alt_2 (R^3)\)?

OpenStudy (experimentx):

it is define that way ... all you need is to check weather W(x1, x2) is bilinear or not.

OpenStudy (experimentx):

\( Alt_2 (\Bbb R^3) \) is probably your functional space.

OpenStudy (loser66):

It says K \(\in Alt_2\) does it imply that there is \(Alt_1\)?

OpenStudy (experimentx):

no .. Alt_1 is no longer bilinear form. for bilinear you need two elements. Alt_1 is your regular linear map.

OpenStudy (loser66):

If I read the solution, I can (accompanied with your support!!:)) understand it. However, if I don't have the solution, I don't know how to construct the function K

OpenStudy (loser66):

Trivially, I have just standard basis of R^3, say x1,x2,x3 how about y1,y2,y3? where do they come from? from dual space of R^3?

OpenStudy (experimentx):

define\( K: \Bbb R^3 \times \Bbb R^3 \to \Bbb R \) by $$K(w_1, w_2) = \det (w_1, w_2, (1,1,0)^T)$$ determinant of this thing. one extra element is added to it. keeping one, you can show the map is linear for other element. and probably Alt_1 does not exist. IDK http://en.wikipedia.org/wiki/Alternatization

OpenStudy (loser66):

Is there any way else to defined K? It is always = det ( something depends on the given condition?)

OpenStudy (kinggeorge):

The idea that I have, is to try and modify the cross product. The cross product is bilinear and alternating, so it should be possible to make a slight alteration and get what we want. For example, we could define for two vectors \(x=(x_1,x_2,x_3),y=(y_1,y_2,y_3)\) the map\[x\oplus y=a_1\vec{i}\]where \(x\times y=a_1\vec{i}+a_2\vec{j}+a_3\vec{k}\) is the traditional cross product. I think this will be a bilinear alternating form that is not identically 0.

OpenStudy (experimentx):

well ... it seems it works for R^3.

OpenStudy (loser66):

so? a1= x2y3-x3y2 a2= x1y3-x3y1 a3=x1y2-x2y1 then?

OpenStudy (experimentx):

SxS is always zero

OpenStudy (kinggeorge):

You shouldn't need to prove anything down from the dirty definitions like that. Alternating and bilinearity follow from the fact that the cross product itself is alternating and bilinear. It's non-zero because \(\vec{j}\oplus\vec{k}=\vec{i}\). However, \(\vec{i}\oplus\vec{j}=\vec0\).

OpenStudy (loser66):

\(\bigotimes\) or direct sum?

OpenStudy (kinggeorge):

I'm just using \(\oplus\) as a symbol for an operation. Also, there's a slight error in my original definition. If we want this to be a form, it needs to map vectors in \(\mathbb{R}^3\) to \(\mathbb{R}\), and not more vectors. So instead of taking \(a_1\vec{i}\), just take the coefficient \(a_1\).

OpenStudy (kinggeorge):

If you don't like \(\oplus\), you can use \(*\) or \(\otimes\) or \(\cdot\), etc.

OpenStudy (experimentx):

I am not so sure about the multilinear alternating form ... but for your particular case, according to your pdf file, ,,, if u and v be two linearly independent vectors in R^3, just take another vector in the subspace of these two vectors. and define the map by it's determinant.

OpenStudy (loser66):

can we have K : R^3 to R^2? is it alternating still?

OpenStudy (kinggeorge):

You mean from \(\mathbb{R}^3\times\mathbb{R}^3\to\mathbb{R}^2\)? I don't think you can have a form that does that since forms need to map to fields.

OpenStudy (loser66):

@experimentx I have to prove the map K is alternating one AND, k-linear form on 3-dimension (the text book said that) So, if I defined K like what you say, the very first step is proving those conditions

OpenStudy (loser66):

@KingGeorge so, nothing to do with the basis? we don't have to translate x = b1x1+b2x2 +b3x3, and y = c1x1+c2x2+c3x3, then define K(x,y) =K(b1x1+b2x2 +b3x3, y) = b1K(x1,y) +...........

OpenStudy (loser66):

x1,x2,x3, is standard basis on R^3, any vector on R^3 can be represented by linear combination of them. so that I have x , y like what I did above

OpenStudy (loser66):

I go from definition of k -linear form and construct K (b1x1+b2x2+b3x3, c1y1+c2y2+c3y3) = b1c1K(x1,x1) +b1c2K(x1,x2) +b1c3K(x1,x3)+b2c1K(x2,x1)+..... and any K(x_i,xj)=0 (given condition) so that K (x,y) = b1c1K(x1,x1) +b2c2K(x2,x2) +b3c3K(x3,x3) and stttttttttuuuuuuck.

OpenStudy (kinggeorge):

To prove the map that I gave is alternating and bilinear, let \(\vec{x},\vec{y},\vec{z}\in\mathbb{R^3}\) and \(c\in\mathbb{R}\). Then let\[\vec{x}\times\vec{y}=a_1\vec{i}+a_2\vec{j}+a_3\vec{k}\]Then\[(c\vec{x})\times\vec{y}=c(\vec{x}\times\vec{y})=c(a_1\vec{i}+a_2\vec{j}+a_3\vec{k}),\]so \((c\vec{x})\oplus\vec{y}=ca_1=c(\vec{x}\oplus\vec{y})\). By bilinearity of the cross product, this means that \(\vec{x}\oplus(c\vec{y})=c(\vec{x}\oplus\vec{y})\) as well.

OpenStudy (kinggeorge):

Also, if\[\vec{z}\times\vec{y}=b_1\vec{i}+b_2\vec{j}+b_3\vec{k}\]then\[(\vec{x}+\vec{z})\times\vec{y}=\vec{x}\times\vec{y}+\vec{z}\times\vec{y}=(a_1+b_1)\vec{i}+(a_2+b_2)\vec{j}+(a_3+b_3)\vec{k}.\]So \[(\vec{x}+\vec{z})\oplus\vec{y}=a_1+b_1=\vec{x}\oplus\vec{y}+\vec{z}\oplus\vec{y}.\]Again since the cross product is bilinear we also have that\[\vec{x}\oplus(\vec{y}+\vec{z})=\vec{x}\oplus\vec{y}+\vec{x}\oplus\vec{z}.\]So this new product is bilinear.

OpenStudy (kinggeorge):

To show it's alternating, we know that \[-\vec{x}\times\vec{y}=\vec{y}\times\vec{x}=-a_1\vec{i}-a_2\vec{j}-a_3\vec{k}\]So \[\vec{y}\oplus\vec{x}=-a_1=-\vec{x}\oplus\vec{y}.\]So the map is also alternating. Thus, it's an alternating bilinear form. Since \(\vec{j}\oplus\vec{k}=1\) it's not the zero form, but \(\vec{i},\vec{j}\) are two linearly independent vectors and \(\vec{i}\oplus\vec{j}=0\). So it satisfies what you need to it.

OpenStudy (loser66):

I got it. Thank you so much @KingGeorge and @experiment . Completely understood. :)

OpenStudy (loser66):

@experimentx hihihi.not @experiment

OpenStudy (experimentx):

lol ... I did nothing more than to create confusion lol.

OpenStudy (loser66):

I can fly now, hehehe...

OpenStudy (kinggeorge):

Thinking of examples for these kinds of things is something I definitely need practice on. I can do the theory, but not the examples...

OpenStudy (loser66):

I still have other problem , I will post and need you guys' brains :) Please

OpenStudy (experimentx):

it seems that in what i wrote above " you need to just take another vector OTHER than in the subspace of these two vectors. and define the map by it's determinant." Determinant is alternating and multilinear map. Helps me too. haha B(u, v) = -B(v, u) ad B(u, u) = 0.

OpenStudy (experimentx):

gotta go too ... see you around.